OpenStudy (anonymous):

Write an equation to model each situation. Then solve: Suppose a video store charges nonmemebers $4 to rent each video. A store membership costs $21 and members pay only $2.50 to rent each video. For what number of videos is the cost the same?

OpenStudy (osanseviero):

2) Same concept nonmembers: c=4m members: c=21+2.5m 4m=21+2.5m 1.5m=21 m=14 (14 movies) FULL THING: http://answers.yahoo.com/question/index?qid=20100920151448AA0deFn